# The Feasibility of Topic-Based Watermarking on Academic Peer Reviews

### v1

@@ 4 Experiments
To evaluate the applicability of topic-based watermarking (TBW) in the domain of peer review, we conduct a series of experiments across multiple dimensions, including text quality, robustness to paraphrasing, and classifier-based attribution.
@@ 4.1 Generation Quality
To assess the impact of TBW on peer review generation, we evaluate outputs using perplexity and BERTScore Zhang et al. (2019). Following prior work Nemecek et al. (2024), we apply a semantic similarity threshold of $\tau=0.7$ to construct topic-aligned green lists. We use 1,000 samples per model configuration (base, few-shot, fine-tuned), each consisting of approximately $200\pm 5$ tokens. Additional results for a lower threshold ($\tau=0.3$) and comparisons to baseline watermarking schemes are provided in Appendix B.
4.1.1 Perplexity
We compute perplexity using the same model that generated the text (Llama-3.1-8B), with lower values indicating higher fluency. Values above 20 are truncated in visualizations for readability (Figure 1), and the number of retained samples is shown in Table 1. This setup reflects how confidently the model assigns probability to its own output, serving as a proxy for fluency.
Figure 1: Perplexity distributions across model configurations with and without TBW ($\tau=0.7$). Lower values indicate better fluency. Values above 20 are truncated for clarity.
@@ 4.1.1 Perplexity
Few-shot NW 1000
TBW 1000
Fine-tuned NW 1000
TBW 1000
Table 1: Number of retained generations with perplexity $\leq$ 20, comparing no watermark (NW) and TBW across model configurations.
TBW introduces only a slight increase in perplexity, consistent with prior findings Nemecek et al. (2024). In the base model, over 50% of unwatermarked generations exceed a perplexity of 20, while nearly all TBW outputs fall below this threshold. This suggests that TBW preserves naturalness and may even enhance lexical consistency in low-context settings by nudging generation toward topic-relevant vocabulary.
4.1.2 BERTScore Evaluation
We use BERTScore F1 to evaluate semantic similarity between generated reviews and ground-truth references. This metric, which compares contextual embeddings, is tolerant to paraphrasing and thus well-suited for open-ended review generation. Results across all model configurations are shown in Figure 2.
Model Attack Setting ROC-AUC Best F1 Score TPR@1%FPR TPR@10%FPR
Base No Attack 0.9678 0.9546 0.9080 0.9560
PEGASUS 0.9359 0.8928 0.7460 0.8610
DIPPER 0.9221 0.8568 0.6690 0.8260
Few-shot No Attack 0.7286 0.7677 0.6260 0.6690
PEGASUS 0.7221 0.7584 0.6090 0.6550
DIPPER 0.7647 0.7537 0.5650 0.6590
Fine-tuned No Attack 0.9813 0.9266 0.8170 0.9480
PEGASUS 0.9435 0.8584 0.5930 0.8260
DIPPER 0.9064 0.8605 0.3480 0.5980
@@ 4.2 Robustness to Paraphrasing Attacks
We assess TBW’s resilience to paraphrasing attacks, a realistic threat model wherein reviewers may rephrase LLM-generated reviews to evade detection while preserving meaning. We focus on full-paraphrase attacks, which best reflect plausible reviewer behavior, and exclude token-level or partial edits.
To align with prior experiments, we generate 1,000 samples per model (base, few-shot, fine-tuned), each with $\sim$200 tokens, using $\tau=0.7$ for topic alignment. Paraphrasing is applied using PEGASUS and DIPPER, the latter configured with lexical = 60 and order = 40, following standard robustness benchmarks Hou et al. (2024); Liu and Bu (2024).
Detection uses the TBW statistical test (see Section 3.2.3), applied to both original and paraphrased generations. Table 2 reports accuracy under three conditions: no paraphrasing, PEGASUS, and DIPPER.
TBW maintains strong robustness in base and fine-tuned models across attack types. The few-shot configuration, however, shows reduced recall (0.6260 $\rightarrow$ 0.5650 under DIPPER), likely due to topic mismatch between prompt examples and the target paper, which weakens topic alignment and reduces detectability post-paraphrasing.
Finally, we verify TBW does not yield false positives on human-written reviews, owing to a partitioning strategy that preserves vocabulary diversity across green lists. For full ROC curves and comparisons with baseline watermarking schemes, see Appendix C.
@@
Classifier Model Watermark Accuracy Precision Recall F1
BERT Base NW 0.290 0.353 0.328 0.278
TBW 0.321 0.346 0.342 0.317
Few-shot NW 0.403 0.373 0.379 0.360
TBW 0.437 0.366 0.369 0.358
Fine-tuned NW 0.400 0.367 0.370 0.364
TBW 0.416 0.366 0.367 0.366
RoBERTa Base NW 0.486 0.344 0.341 0.305
TBW 0.432 0.357 0.352 0.350
Few-shot NW 0.399 0.362 0.368 0.337
TBW 0.424 0.371 0.371 0.353
Fine-tuned NW 0.406 0.367 0.374 0.367
TBW 0.443 0.401 0.403 0.402
4.3 Classifier-Based Attribution
To complement watermark detection, we evaluate whether LLM-generated peer reviews can be attributed to their original review labels (e.g., accept, borderline, reject) using standard classification models. This task provides a content-based signal of semantic alignment, helping assess whether watermarking affects the interpretability or label consistency of generated reviews. We frame this as a three-way classification problem based on the review score originally assigned to each paper.
4.3.1 Data and Training Protocol
We first construct a labeled dataset by extracting review texts from our generation pipeline and assigning a class label based on the associated ground truth rating (e.g., scores 1–4 mapped to reject, 5–6 to borderline, and 7–10 to accept). To ensure accurate mapping, we align generated reviews with their original metadata using paper titles as unique identifiers. The final dataset consists of generated reviews paired with class labels, drawn from the fine-tuned generation split described in Section 3.1.4.
We train two transformer-based classifiers, BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019), to predict the rating category of each review. The dataset is stratified into training and held-out test splits, with 9,000 balanced training samples (3,000 per class) and 1,000 test samples. Tokenization is performed using each model’s native tokenizer, and models are fine-tuned using the HuggingFace Trainer API with early stopping based on F1. We adopt 4-bit precision, label smoothing (0.1), and a cosine learning rate schedule with warmup. Additional training hyperparameters and evaluation on the testing set are provided in Appendix D.
4.3.2 Evaluation
Once trained, both classifiers are applied to a held-out set of generated reviews produced by various generation configurations (base, few-shot, fine-tuned) with and without TBW using $\tau=0.7$. For each review, we extract the title from the input prompt, retrieve the associated ground truth score from metadata, and map it to a label for evaluation. We evaluate model performance with and without TBW to assess whether watermarking impairs label recoverability.
As shown in Table 3, we observe no degradation in classification performance due to TBW. On the contrary, in most configurations, applying TBW leads to modest improvements in both accuracy and F1. This suggests that topic-based watermarking preserves the semantic structure necessary for accurate label prediction and even enhance it by encouraging more topically consistent language. These findings reinforce TBW’s suitability for attribution tasks in domain-sensitive contexts like peer review, where both traceability and semantic fidelity are critical. Additional analysis on class-specifics for human-written reviews is provided in Appendix E and classifier attribution performance under a lower topic similarity threshold ($\tau$ = 0.3) to assess the impact of weaker topic alignment is provided in Appendix F. For an analysis of how review content and structure shift under paraphrasing, see Appendix G, which provides changes in accuracy under paraphrasing.
5 Discussion
@@
From a deployment perspective, TBW offers a practical solution for reviewer attribution. The method is efficient and detection incurs minimal computational overhead, making it suitable for integration into existing conference submission pipelines Nemecek et al. (2024). Its low latency and lack of architectural modifications make it a compelling candidate for enforcement mechanisms in venues that prohibit LLM-assisted review writing.
@@
6 Conclusion
We present a comprehensive evaluation of topic-based watermarking in the context of academic peer review, a high-stakes domain where LLM use is often restricted but difficult to detect. Unlike prior work that focuses on general-purpose text, our study demonstrates that topic-based watermarking can preserve generation quality, maintain robustness under paraphrasing, and support attribution across different LLM configurations. Its semantic grounding and low computational overhead make it a practical solution for enforcing LLM usage policies in peer review, offering a minimally intrusive mechanism to help safeguard the integrity of academic evaluation.
Limitations
This work inherits a key limitation of topic-based watermarking: the topic-matching assumption. As noted in the original proposal Nemecek et al. (2024), watermark detection may degrade if the semantic topic of the generated output drifts significantly from the original prompt. This is particularly challenging in open-domain generation, where the input prompt is often unavailable at detection time. However, in the context of peer review, this limitation is largely mitigated. Reviewers must prompt the LLM using the content of the paper, either by directly including the text or referencing its abstract and title, ensuring that the generated review remains topically aligned with the source. Furthermore, during detection, conference organizers have access to the submission itself, allowing them to reliably identify the intended topic and recover the correct green list. As a result, the topic-matching assumption holds in this use case.
Ethical Considerations
Appendix A Peer Review Task Specifics
This appendix provides additional details regarding the peer review generation setup described in Section 3.1. Specifically, we include conference-level review statistics and implementation details for fine-tuning the Llama-3.1-8B model.
A.1 Conference Review Statistics
@@
A.2 Fine-tuning Details

### v2

@@ 4 Experiments
We conduct a series of experiments across multiple dimensions, including text quality, robustness to paraphrasing, and classifier-based attribution.
@@ 4.1 Generation Quality
To assess TBW’s impact on generation quality, we evaluate perplexity across 1,000 samples per model configuration (base, few-shot, fine-tuned), comparing against unwatermarked baselines and two existing schemes: KGW kirchenbauer2023watermark and Google’s SynthID-Text dathathri2024scalable. Each sample contains approximately $200\pm 5$ tokens. Parameter configurations, additional threshold results, human evaluations, and BERTScore evaluations are detailed in §D.
We compute perplexity using the generating model (Llama-3.1-8B) as a fluency proxy, where lower values indicate higher naturalness. For visualization clarity, values above 20 are truncated in Figure 1.
Figure 1: Perplexity distributions across model and watermark configurations. Lower values indicate better fluency. Values above 20 are truncated for clarity.
The results demonstrate that TBW introduces minimal perplexity degradation while maintaining fluency compared to existing watermarking schemes. While Figure 1 shows TBW with slightly higher median perplexity than KGW in some configurations, the truncation at perplexity 20 obscures an important quality difference. Table 1 reports the number of samples retained below this threshold across all watermarking schemes and model configurations.
@@
KGW 840
SynthID 538
Few-shot All schemes 1000
Fine-tuned All schemes 1000
Table 1: Number of retained generations with perplexity $\leq 20$ across model configurations. TBW retains 991/1000 base model samples, significantly outperforming KGW (840/1000) and SynthID (538/1000), demonstrating superior outlier suppression.
Notably, TBW retained 991 of 1000 base model samples below the perplexity threshold, excluding only 9 outliers, whereas KGW retained 840 (160 excluded) and SynthID retained only 538 (462 excluded). This demonstrates that TBW produces far fewer low-quality outliers than existing methods, the best performance among all watermarking schemes tested. These results suggest TBW preserves generation naturalness but enhances lexical consistency in low-context scenarios by steering generation toward topic-relevant vocabulary.
While perplexity serves as a useful fluency proxy, it does not fully capture semantic fidelity to the original content. To complement this analysis, we evaluate semantic similarity using BERTScore zhang2019bertscore, which measures contextual embedding alignment between generated reviews and ground-truth references (§D.2), and conduct a small-scale human evaluation on the fluency, coherence, and usefulness of generated reviews (§D.3). Furthermore, we evaluate TBW under a relaxed similarity threshold ($\tau=0.3$) to assess sensitivity to green-list construction (§D.4). These complementary metrics confirm that TBW preserves both fluency and semantic integrity across generation conditions.
@@ 4.2 Robustness to Paraphrasing Attacks
We evaluate TBW’s resilience against paraphrasing attacks, which represent a realistic threat model where reviewers may rephrase LLM-generated content to evade detection while preserving semantic meaning. We focus on full-text paraphrasing attacks that best reflect plausible reviewer behavior, excluding token-level or partial edit scenarios.
Detection Accurcy
Model Attacks TBW KGW SynthID
Base No Attack 0.946 0.971 0.909
PEGASUS 0.847 0.477 0.135
DIPPER 0.876 0.754 0.173
Few-shot No Attack 0.622 0.975 0.959
PEGASUS 0.580 0.580 0.359
DIPPER 0.517 0.748 0.225
Fine-tuned No Attack 0.880 0.926 0.960
PEGASUS 0.583 0.437 0.180
DIPPER 0.584 0.657 0.159
Table 2: Detection accuracy of TBW, KGW, and SynthID across model configurations and paraphrasing attack types. Scores reflect the proportion of correctly identified watermarked samples per condition.
We generate 1,000 samples per model configuration (base, few-shot, fine-tuned) for each watermarking scheme, with each sample containing $200\pm 5$ tokens. Paraphrasing attacks are implemented using two established models: PEGASUS zhang2020pegasus and DIPPER krishna2023paraphrasing (configured with lexical=60 and order=40), following standard robustness evaluation protocols hou-etal-2024-semstamp; liu2024adaptive.
Table 2 shows detection accuracy across three attack conditions. TBW demonstrates superior robustness compared to existing methods, maintaining detection accuracy for base models and fine-tuned models even under paraphrasing attacks. In contrast, KGW and SynthID show degradation, with SynthID dropping to as low as 13.5% accuracy under PEGASUS attacks.
The few-shot configuration exhibits reduced performance across all methods, with TBW achieving 58% accuracy under PEGASUS and 52% under DIPPER. This degradation likely stems from topic misalignment between the prompt examples and target papers, which weakens topic alignment and reduces detectability post-paraphrasing.
Importantly, this represents a characteristic trade-off of TBW’s semantic design. While topic-aware watermarking provides robustness in controlled settings (base and fine-tuned models), it becomes sensitive to exemplar-target mismatch in few-shot prompting. This sensitivity surfaces an important consideration for effective deployment of TBW in few-shot scenarios, which would benefit from improved exemplar selection strategies or dynamic prompt construction that ensures topical consistency between examples and the target generation task.
We verify that TBW maintains low false positive rates on human-written reviews through its vocabulary partitioning strategy, which preserves lexical diversity across topic-specific green lists. Complete ROC analysis is provided in §E.
4.3 Classifier-Based Attribution
Beyond detecting the presence of watermarks, one question is whether watermarking degrades the semantic integrity of generated reviews in ways that affect downstream interpretation tasks. Specifically, if watermarking distorts the linguistic features that correlate with review sentiment (accept, borderline, reject), it could undermine the very purpose of generating coherent and evaluable reviews.
To address this concern, we evaluate whether both watermarked and non-watermarked LLM-generated peer reviews can be accurately attributed to their original review labels (accept, borderline, reject) using transformer-based classification models. This experiment serves two purposes: (i) it validates that watermarked reviews preserve sufficient semantic structure to support rating prediction, and (ii) it tests whether TBW’s topic-aware biasing introduces artifacts that either help or hinder sentiment classification. If watermarking significantly degrades classifier performance, it would suggest that the semantic alterations required for attribution compromise review quality in measurable ways.
Base Few-shot Fine-tuned
Metric NW TBW NW TBW NW TBW
BERT
Accuracy 0.290 0.321 0.403 0.437 0.400 0.416
Precision 0.353 0.346 0.373 0.366 0.367 0.366
Recall 0.328 0.342 0.379 0.369 0.370 0.367
F1 0.278 0.317 0.360 0.358 0.364 0.366
RoBERTa
Accuracy 0.486 0.432 0.399 0.424 0.406 0.443
Precision 0.344 0.357 0.362 0.371 0.367 0.401
Recall 0.341 0.352 0.368 0.371 0.374 0.403
F1 0.305 0.350 0.337 0.353 0.367 0.402
@@ 4.3 Classifier-Based Attribution
We use a balanced dataset of 9,000 training samples (3,000 per class) and 1,000 test samples. Reviews are mapped to three categories based on original scores: 1-4 (reject), 5-6 (borderline), and 7-10 (accept). We train BERT devlin-etal-2019-bert and RoBERTa zhuang-etal-2021-robustly classifiers to predict review ratings based on generated review text, evaluating performance on both watermarked and non-watermarked reviews.
As shown in Table 3, TBW demonstrates mixed but generally positive effects on classification performance. While some configurations show modest degradation (e.g., RoBERTa base accuracy), the majority of results indicate that TBW causes little to no degradation and often leads to improvements in both accuracy and F1 scores. We hypothesize that this occurs because topic-based watermarking encourages more topically consistent language that aligns better with the underlying review content.
These findings reinforce TBW’s suitability for domain-sensitive contexts like peer review, where both traceability and semantic fidelity are critical. The results demonstrate that watermarking does not compromise and may even enhance the interpretability of generated reviews with respect to their underlying evaluative stance. Additional experimental details, including class-specific performance metrics, evaluation under lower similarity thresholds ($\tau=0.3$), and robustness evaluation under paraphrasing attacks, are provided in §F.
5 Conclusion
We present the first systematic evaluation of topic-based watermarking (TBW) for academic peer review, demonstrating that TBW preserves generation quality while maintaining robust detection under paraphrasing attacks. Our experiments across base, few-shot, and fine-tuned LLM configurations show that TBW’s semantic grounding naturally aligns with peer review constraints, where content must remain topically consistent with the evaluated paper. These findings highlight TBW’s viability as a minimally intrusive solution for LLM attribution in peer review settings, offering a practical mechanism to safeguard academic evaluation integrity. Discussion of deployment considerations and limitations is provided in §B.
Limitations
This work inherits a key limitation of topic-based watermarking: the topic-matching assumption. As noted in the original proposal nemecek2024topic, watermark detection may degrade if the semantic topic of the generated output drifts significantly from the original prompt. This is particularly challenging in open-domain generation, where the input prompt is often unavailable at detection time. However, in the context of peer review, this limitation is largely mitigated. Reviewers must prompt the LLM using the content of the paper, either by directly including the text or referencing its abstract and title, ensuring that the generated review remains topically aligned with the source. Furthermore, during detection, conference organizers have access to the submission itself, allowing them to reliably identify the intended topic and recover the correct green list. As a result, the topic-matching assumption holds in this use case.
Thirdly, our evaluation of TBW’s robustness in the context of academic peer reviews focuses on full-text paraphrasing attacks, which represent strong adversarial models where the entire review is rephrased. We selected this threat model because it represents the most comprehensive transformation: if TBW remains detectable when all text is paraphrased, it should naturally be robust to partial edits affecting only portions of the review. However, we acknowledge that real-world reviewer behavior may include additional manipulation strategies not evaluated in this work, such as hybrid human-LLM composition (mixing original writing with generated content), section reordering, selective compression or expansion of specific review components, and multi-stage translation attacks (e.g., English to an intermediate language and then back to English). While these scenarios represent valuable directions for future robustness analysis, we expect TBW’s semantic grounding to provide resilience against such attacks, particularly in cases where the topical alignment between the paper and review is preserved. Future work should systematically evaluate these additional threat models to fully characterize TBW’s operational boundaries.
An important distinction not addressed in our current evaluation is differentiating between text originally authored by a human reviewer but subsequently polished by an LLM (e.g., grammar correction, style refinement) versus text entirely generated by an LLM. This represents a challenging but critical use case, as many conferences permit LLM-assisted writing improvement while prohibiting wholesale generation. TBW’s detection mechanism, which relies on green-list token density, would likely flag both scenarios similarly since the polishing process may still introduce watermarked tokens. Developing attribution methods that can reliably distinguish between these two cases remains an important direction for future work and may require complementary techniques beyond watermarking alone.
Ethical Considerations
Acknowledgments
This work was partially supported by the National Science Foundation (NSF) under grant numbers 2050410 and OAC-2112606. ChatGPT and Claude were used in the preparation of this work for editing and grammar checking. No passages were copied without full author review, and all substantive ideas, analysis, and conclusions are the product and responsibility of the authors. Additionally, the listed AI tools were utilized for code development and early manuscript reviews.
Appendix A Related Work
Since the release of ChatGPT, LLMs have been rapidly adopted across various stages of the academic workflow. Their use has raised concerns about authorship and peer review integrity. Most conferences and journals now permit authors to leverage LLMs; however, this permissive stance does not extend to peer reviewers. Leading venues such as NeurIPS and ACL explicitly prohibit the use of LLMs by reviewers neurips2025llm; emnlp2025review. These policies reflect growing concerns around review quality, including the risk of shallow or hallucinated feedback, reduced technical depth, and breaches of confidentiality that would compromise the double-blind review process li2024use.
Despite these restrictions, recent studies suggest that LLM-assisted reviews are already present at major conferences. liang2024monitoring estimate that 5-15% of reviews were substantially modified using LLMs, with affected reviewers showing lower confidence and less engagement during rebuttals. latona2024ai report similar trends and observe a score inflation effect, while ye2024we show that even subtle LLM manipulations can shift paper rankings. Together, these findings underscore the risks unauthorized LLM use poses to peer review fairness and rigor.
Given the increasing use of LLMs for peer review generation, recent work has focused on detecting and attributing such content. Much of this research explores classifier-based detection or semantic similarity methods aimed at identifying AI-generated text. For example, yu2025your propose a detection method based on the semantic similarity between a known LLM-generated review and a test review, flagging a review as machine-generated when similarity exceeds a threshold. Similarly, kumar2025mixrevdetect introduce a partition-based method under the assumption that a review contains both human- and LLM-written components. They segment the review into distinct points, complete each segment with a reference LLM, and measure semantic similarity between these completions and the original text to detect potential LLM involvement.
However, these detection methods fail under paraphrasing or hybrid-review scenarios, where even minor edits or partial human rewriting can evade detection. To address this limitation, watermarking offers a promising alternative by embedding identifiable signals directly into the generated text. One foundational method is the KGW algorithm kirchenbauer2023watermark, which partitions the model’s vocabulary into “green” and “red” token sets. During generation, the model is subtly biased to sample more frequently from the “green” list, which acts as a watermark-carrying set, while avoiding tokens in the “red” list. This results in output text that biases outputs toward “green” tokens with minimal quality loss. Variants aim to improve robustness and preserve quality liu2024a; zhao2023provable; hou-etal-2024-semstamp.
More recently, commercial systems have also entered this space. For example, Google’s SynthID-Text watermarking system employs a strategy called Tournament Sampling, in which candidate tokens are ranked according to randomized watermarking functions, and the highest-ranked token is selected during generation dathathri2024scalable. While both academic and commercial watermarking approaches have shown promise, they are primarily evaluated on general-purpose domains such as news or encyclopedic text, and rarely tested under the stylistic and ethical constraints found in peer review.
While a few frameworks target peer review watermarking rao2025detecting, they rely on tightly integrated pipelines and lack evaluation across adaptation modes. Topic-based watermarking (TBW) nemecek2024topic, originally proposed for open-domain text, provides a lightweight, semantically guided alternative.
Appendix B Discussion
@@
From a deployment perspective, TBW offers a practical solution for reviewer attribution. The method is efficient and detection incurs minimal computational overhead, making it suitable for integration into existing conference submission pipelines nemecek2024topic. Its low latency and lack of architectural modifications make it a compelling candidate for enforcement mechanisms in venues that prohibit LLM-assisted review writing.
@@
Appendix C Peer Review Task Specifics
This appendix provides additional details regarding the peer review generation setup described in §3. Specifically, we include conference-level review statistics and implementation details for TBW as well as prompting and fine-tuning the Llama-3.1-8B model.
C.1 Conference Review Statistics
@@
C.2 TBW Parameter Details
C.2.1 Implementation Components
We use the same core components as the original TBW implementation to ensure consistency. Token and topic embeddings are computed using the all-MiniLM-L6-v2 sentence embedding model reimers-2020-multilingual-sentence-bert. Topic extraction from input prompts is performed using KeyBERT grootendorst2020keybert for keyword-based topic identification.
C.2.2 Topic Assignment & Green List Construction
Following the TBW framework, we partition the vocabulary into green lists based on semantic similarity to our predefined set of $K=4$ domain-specific topics: {theory, applications, models, optimization}. Each token $v\in V$ is assigned to the green list $G_{t_{i}}$ if its cosine similarity with topic $t_{i}$ exceeds threshold $\tau$:
$\text{sim}(v,t_{i})=\frac{\mathbf{e}_{v}\cdot\mathbf{e}_{t_{i}}}{\|\mathbf{e}_{v}\|\|\mathbf{e}_{t_{i}}\|}$
Tokens not meeting the similarity threshold are placed in a residual set and evenly distributed across all green lists to maintain complete vocabulary coverage.
C.2.3 Topic Granularity Selection
We selected $K=4$ domain-specific topics to balance semantic coverage with watermark detection robustness. Finer-grained topic divisions (larger $K$) would create sparser green lists $G_{t_{i}}$, potentially weakening watermark signals and reducing robustness to paraphrasing. Coarser divisions (smaller $K$) would sacrifice semantic alignment between prompts and outputs. The selected topics: {theory, applications, models, optimization}, capture the major thematic dimensions of ML conference reviews while maintaining sufficient green-list density for reliable detection. Future work could explore adaptive topic granularity strategies, such as hierarchical topic structures or automatic topic discovery from review corpora.
C.2.4 Generation & Detection Parameters
We apply a logit bias of $\delta=2.0$ to green-list tokens during generation, consistent with values reported in prior watermarking literature kirchenbauer2023watermark. For token-to-topic assignment, we primarily use a cosine similarity threshold of $\tau=0.7$, with additional evaluation at $\tau=0.3$ to assess how watermark detection and text quality vary under relaxed alignment constraints.
For detection, we use the statistical z-test with baseline proportion $\gamma$ set to the expected green-token rate under random sampling, and threshold $z_{\text{threshold}}$ tuned for desired false positive rates.
C.3 Model Configuration Details
C.3.1 Few-shot Prompting Setup
Title: [TITLE] Abstract: [ABSTRACT] Please write a detailed review.
C.3.2 Fine-Tuning Setup
For fine-tuning, we follow a supervised instruction-tuning setup where each instance consists of an input prompt (title + abstract + instruction) and a target completion (review text). The dataset is split into training (80%), validation (10%), and test (10%) subsets. We fine-tune using LoRA (Low-Rank Adaptation) with 4-bit quantization, enabling gradient checkpointing and early stopping. The objective is to improve the fluency and consistency of generated reviews while approximating the tone and structure typical of human-written peer reviews.